graceyhansen09:

Bella wants to write two equations to model the streets on this map. She can use y = x – 4 to describe Platte Way. Find one absolute value equation to describe Marteen Rd and Smith St.

umm:

Answer choices: y = | x - 2 | - 3 y = | x + 2 | - 3 y = | x + 3 | - 2 y = | x - 3 | - 2